# algorithmic

### v2

@@ 3.1 Attributes of Feedback Types
Feedback types may differ on many dimensions, some of which relate to the way feedback is given (arity, involvement), others to the form of the query instance it is given on (granularity, abstraction), and yet others to features of the human interaction (intent, explicitness). The attributes we discuss in this section are based on the framework proposed by Metz et al. (2023). We have adjusted and added terminology where it aids clarity, generalized the distinction between relative and absolute feedback to ‘arity’, and added the categories of co-generative involvement and literal intent. Furthermore, we systematically analyze a set of exemplary feedback types in the next section, expanding on the initial examination of a smaller range of abstract classes by Metz et al. (2023). In the following, we will introduce each of the six attributes in more detail.
@@
This attribute describes whether a single instance is evaluated in isolation (unary) or relative to other instances (binary, $n$-ary). Unary feedback is often convenient for detailed and descriptive feedback but lacks any grounding and therefore puts a great burden on the human to provide consistent feedback. Non-unary feedback always has an implicit grounding but relates to the instances being comparable. While $n$-ary feedback, such as rankings, can provide more information than binary feedback, it also puts a higher cognitive burden on the labeler.
@@
The labeler may either passively observe an instance, actively generate it, or coactively participate in its generation (co-generation). Passive involvement poses the smallest challenge to the labelers since it does not require the ability to demonstrate the task. It can also easily be directed at the most informative examples with active learning techniques. Unfortunately, passive feedback often cannot match the information density of active feedback. It is, therefore, common to combine both types to first initialize the reward model from (possibly very suboptimal) active feedback and then refine it from passive feedback. Between these two extremes is co-generative feedback, in which a human can share control with the agent. This can be less demanding than active feedback and makes it possible to direct the human’s attention to the most informative samples, but it is still more taxing than purely passive involvement.
@@
Feedback may also differ on the granularity of the instances being evaluated. This ranges from whole episode recordings over partial segments to feedback on individual steps (i.e., states, actions, or state-action pairs). A more coarse-grained granularity has the advantage of giving humans more context and getting feedback for larger sections of behavior but also poses credit assignment problems. Finer-grained feedback is much easier to learn from and simplifies credit assignment, but is often impractical or tedious for humans to provide. It is also possible to strike a compromise, as demonstrated by Guan et al. (2021) who propose to use queries on step granularity, but batch queries within an episode to provide additional context and make it easier to give feedback. Note that we only classify a type of feedback as “episode” granularity if it requires entire episodes. If it is compatible with partial segments as well, we classify it as “segment” even if the discussed source paper uses entire episodes.
@@
This describes whether feedback is given directly on raw instances, e.g., behavior recordings (see granularity) or on abstract features of the instances. While feature-level information can be easier to learn from, extracting useful features is challenging. In some contexts, it may also be harder for a human to make abstract judgments rather than more intuitive instance-level judgments. Note that this always refers to the level of abstraction that the user sees, which may differ from the features used as inputs for the reward model. Types of feedback that depend on active generation (see involvement), such as improvements, generally work on a raw instance level.
@@
Humans may communicate explicitly for the purposes of feedback or implicitly as a side-effect of actions directed at other purposes. While explicit information is easiest to learn from, implicit information is often much more readily available and can possibly communicate more detailed preferences.
@@
The distinction between literal and pedagogical feedback was introduced by Milli & Dragan (2020). They argue that humans with pedagogical intent (i.e., evaluative, instructive, or descriptive) may act differently compared to humans with literal intent. Even though they find that assuming the (wrong) literal intent can still lead to better reward inference, it still indicates that it can be important to know this intent to choose the right human model (see Section 5.1.5).
@@ 3.2 Common Classes
Table 4 gives an overview of the classes and their attributes as described in Section 3.1.
@@ 3.2.1 Primary Feedback Classes
Critique is arguably the most direct type of feedback. In this setting, the human expresses their preference by directly critiquing an instance of agent behavior, often in the form of binary feedback. Note that critique, as considered in this survey, is distinct from directly supplying a reward signal since it is given in an asynchronous and indirect manner (see Section 1.3). The defining features of critique are that the human passively observes the behavior (involvement), gives feedback on a single instance (arity), and does so explicitly (explicitness) with an evaluative intent. The feedback may be given for any granularity and on any level of abstraction.
@@
Binary comparisons and rankings are among the most common types of feedback. The defining features of comparisons are that the human passively observes the behavior (involvement), gives relative feedback on multiple instances (arity), and does so explicitly (explicitness) with an evaluative intent. It is most commonly given on a segment (granularity), but other granularities are possible in principle (Cosner et al., 2022). Similarly, comparisons are commonly requested on an instance level (abstraction), but this is not a requirement.
@@
This basic setting has been extended and modified in various ways. To reduce noise in the labels, it is common to extend the binary choice by giving the labelers the option to avoid the hard choice and instead indicate incomparability, uncertainty, or perceived similarity (Holladay et al., 2016). This option is commonly interpreted as “equally preferable”, e.g., as if each trajectory had an equal probability of being preferred in the preference predictor. It is alco common, however, to additionally provide an “incomparable” option which results in simply omitting the query from the data set (Christiano et al., 2017; Ibarz et al., 2018). In contrast to this, Verma et al. (2023) explicitly state that they do not allow for the equally preferred option, arguing that these cases are rare enough not to matter very much. Another line of research suggests that more precision in the expression of pairwise preferences, such as softening the hard binary choice to scalar feedback indicating the strength of a preference (Wilde et al., 2021), can be beneficial for preference learning. Other extensions change the pairwise setting to choices among larger choice sets (Ziegler et al., 2020) or even full rankings (Myers et al., 2021; Brown et al., 2020; Ouyang et al., 2022). Jain et al. (2015) compare different kinds of re-ranking feedback (select first which is better than top, choose best from top-5, choose best from random set). Their study suggests that the first two are roughly equivalent, while the latter is much less informative. Askell et al. (2021) evaluate binary comparisons as well as ranking, but finds that ranking performs better. Ziegler et al. (2020) note that for language tasks, a larger choice set can amortize the time needed for a labeler to get acquainted with the context necessary to understand a query. Basu et al. (2019) propose to use hierarchical queries, i.e., a sequence of pairwise comparisons that build up on each other.
Pairwise comparisons are generally easier to supply than ratings, even when the feedback is given by a foundation model (Wang et al., 2024b). While absolute feedback is generally dependent on a policy serving as an implicit baseline (MacGlashan et al., 2017), the explicit baseline in pairwise comparisons can be more easily controlled. Therefore, comparisons provide many benefits over absolute ratings, such as reduced bias, inconsistencies, and subjectivity (Yannakakis & Martínez, 2015). On the flip-side, comparisons convey relatively little information per label. Tien et al. (2023) study the weaknesses of pairwise-comparison-based reward learning. Since the amount of information provided for each label is small, these models are prone to causal confusion, i.e., misattributing reward to noise and misidentifying the reward function.
@@
One limitation of trajectory comparisons is that they require a set of roughly comparable trajectories. In many real-world environments, starting conditions or even the agent’s current task may vary between episodes. In these cases, it is hard for human labelers to compare trajectories from these episodes, limiting the usefulness of comparison feedback. One way to remedy this limitation is to provide feedback within a single trajectory. Instead of comparing a set of instances with each other as in regular comparative feedback, inter-temporal feedback conveys relative judgments over different states in time within a single instance. The defining features of inter-temporal feedback are that it is given explicitly (explicitness) on segment (granularity) while passively observing (involvement) a single instance (arity) of the agent’s behavior with evaluative intent. It is most commonly given on raw instances, but any level of abstraction is possible in principle. There are two main ways to convey this feedback: Reward sketching and inter-temporal preferences.
Reward sketching, as introduced by Cabi et al. (2020), involves users sketching a visual representation of the reward function over time. This type of feedback, which can be given by sketching a graph with the mouse while watching a behavior recording, provides intuitive, per-timestep reward annotations. Rahtz et al. (2022) also adopted this approach, referring to it as “one of the highest-bandwidth feedback mechanisms currently available”.
@@
Proxy rewards are partial or inaccurate reward functions that convey information about the task the agent is supposed to complete but may not induce optimal behavior. This form of feedback does not generally refer to any particular behavior instance but instead gives global direction for the entire task. However, in line with our selection criteria (Section 1.3), we only consider proxy reward feedback that is interactive and online within the context of one or multiple observations to fill holes in the initial description. The defining features of proxy reward feedback is that the labeler passively observes the agent’s behavior (involvement) and gives feedback explicitly (explicitness) on a feature-level (abstraction) with descriptive intent (intent). Proxy reward feedback may be given with respect to a single or multiple instances (arity), although it generally refers to multiple instances. It is most commonly given on an episode (granularity), but other granularities are possible in principle.
The work by He & Dragan (2021) exemplifies this form of feedback by using proxy reward functions, i.e., preliminary reward functions that might not cover all edge cases, to guide the agent toward learning the actual reward function. Alternatively, Mindermann et al. (2018); Hadfield-Menell et al. (2017b) suggest querying about the reward function. They allow users to choose from a set of understandable, linear proxy rewards or to specify which features are more critical in the linear reward structure. In a related setting, Guan et al. (2023) lets the user specify changes to a current symbolic reward function in the form of changes to target attributes (e.g., walking speed).
@@
Interventions can be considered to be an instance of the shared autonomy setting since the agent and the user share autonomy to reach a common goal. There are two main ways to leverage interventions: One is to learn form the occurrence of an intervention itself that the prior behavior was suboptimal, as done, e.g., by Luo et al. (2024) by directly inferring negative rewards from interventions without learning a reward model, and the other is to learn from the corrected trajectory. The latter setting is studied by Abramson et al. (2022), who ask humans to intercede on agent failure. They use this to collect targeted demonstrations where the agent is the weakest and to identify challenging situations for their evaluations. The gathered data is then used for behavior cloning and reward model training. The fact that a correction occurred is not directly used as feedback. In addition to the above distinction, interventions can come in two main formats: Either by overtaking control from the agent to correct its behavior or by physically correcting the agent’s movements. Losey et al. (2022) proposes to learn a reward model from such physical corrections, a method that was later extended (Losey & O’Malley, 2018) to incorporate uncertainty for active learning and risk-sensitive deployment. Li et al. (2021) further extend this setting to learn from a sequence of correlated physical corrections without needing to wait until the trajectory is completed.
The correction case is closely related to the setting of coactive learning (Shivaswamy & Joachims, 2015), in which a learning system repeatedly proposes a solution which a user may correct to reach a common goal. Jain et al. (2015) treat corrections as a demonstration while Jeon et al. (2020) propose (but do not evaluate) an alternative interpretation of inferring implicit preferences from comparisons by assuming the corrected trajectory is preferred over the original one.
@@
While much work in language for RL focuses on the problem of learning a policy that follows natural-language instructions in a non-interactive setting (Hermann et al., 2017; Nair et al., 2021), we focus on works where language is used as feedback, not for task definition. Note, however, that RLHF can be a useful tool for learning a language-conditioned policy, regardless of whether or not the feedback itself is in the form of natural language. As an example of this, Abramson et al. (2022) use RLHF in an interactive agents setting, where one agent (responsible for solving problems) is controlled by the learned policy and another agent (responsible for setting tasks) is controlled by the human.
@@
Language feedback can also be interpreted by an LLM directly, without learning a reward model. This is demonstrated in the concurrent works by Ma et al. (2024) and Xie et al. (2024) who propose to learn a symbolic reward model in the form of a language-model written piece of (Python) code. They use natural language feedback to improve the reward model based on observations of the agent’s behavior induced by the previous version of the reward model.
@@ 3.2.2 Supplementary Classes
Merge this with representation-specific, highlight distinction direct feedback about preferences vs feedback about representation in section introduction
This section introduces two feedback classes, e-stops and importance, that can be used in conjunction with a primary class to learn a reward model. These classes are not sufficient to learn a reward model on their own, but can supplement a primary feedback type.
@@
Emergency stops (e-stops) (Ghosal et al., 2023) are an active type of feedback. In this type of feedback, the human may intervene with the agent’s behavior by stopping it, i.e., they may choose to stop the agent’s current trajectory at any point. This is closely related to interventions but, in contrast to those, e-stops do not suggest an alternative action. The defining features of e-stops are that the human both passively observes the agent’s behavior (involvement), gives absolute feedback on a single instance (arity) on the instance level (abstraction) and does so implicitly (explicitness) as a side-effect of regular interaction. The intent is literal due to the implicit nature. For the purposes of intervention, the human usually observes the full episode (granularity). Due to the small amount of infrequent information they provide, e-stops should only be considered as a supplementary feedback type.
E-stops are primarily intended to prevent bad behavior and only implicitly convey information about the correct behavior. This interaction and the arising incentives have been formalized in the form of the “off-switch game” by Hadfield-Menell et al. (2017a). Jeon et al. (2020) propose to interpret this as a form of reward-rational feedback, where the ‘off’ choice maps to the trajectory with the robot remaining still after the off switch has been triggered (see Section 5.1.1). Kahn et al. (2021) demonstrate that a robot can learn to navigate using such feedback.
@@
Another form of supplementary feedback may come in the form of importance labels, communicating which parts of the observation are important for the objective. Its defining features are that the importance information itself does not contribute towards generating behavior samples (observed involvement), is of descriptive intent, and is given explicitly (explicitness). Granularity, arity, and abstraction may vary depending on the primary feedback type. Since importance feedback needs a base task with respect to which the importance is defined, it cannot be used on its own but is rather a supplementary type of feedback.

### v3

@@ 3.1 Attributes of Feedback Types
Feedback types may differ in many dimensions, some of which relate to the way feedback is given (arity, involvement), others to the form of the query instance it is given on (granularity, abstraction), and yet others to features of the human interaction (intent, explicitness). The attributes we discuss in this section are based on the framework proposed by Metz et al. (2023). We have adjusted and added terminology where it aids clarity, generalized the distinction between relative and absolute feedback to ‘arity’, and added the categories of co-generative involvement and literal intent. Furthermore, we systematically analyze a set of exemplary feedback types in the next section, expanding on the initial examination of a smaller set of abstract classes by Metz et al. (2023). In the following, we will introduce each of the six attributes in more detail.
@@
This attribute describes whether a single instance is evaluated in isolation (unary) or relative to other instances (binary, $n$-ary). Unary feedback, e.g., an absolute score for an observed behavior, is often convenient for detailed and descriptive feedback but lacks any grounding and therefore puts a great burden on the human to provide consistent feedback. Non-unary feedback always has an implicit grounding but requires the instances to be comparable. While $n$-ary feedback, such as a ranking over multiple observed behaviors, can provide more information than binary feedback, it also puts a higher cognitive burden on the labeler.
@@
The labeler may either passively observe an instance, actively generate it, or coactively participate in its generation (co-generation). Passive involvement, exemplified by an evaluator that observes, but does not intervene in a robot’s behavior, poses the smallest challenge to the labelers since it does not require the ability to demonstrate the task. It can also easily be directed at the most informative examples with active learning techniques. Unfortunately, passive feedback often cannot match the information density of active feedback such as a human steering a robot to generate a demonstration. It is, therefore, common to combine both types to first initialize the reward model from (possibly very suboptimal) active feedback and then refine it from passive feedback. Between these two extremes is co-generative feedback, in which a human can share control with the agent. This can be less demanding than active feedback and makes it possible to direct the human’s attention to the most informative samples, but it is still more taxing than purely passive involvement.
@@
Feedback may also differ on the granularity of the instances being evaluated. This ranges from feedback over whole episodes over partial segments to feedback on individual steps (i.e., states, actions, or state-action pairs). A more coarse-grained granularity has the advantage of giving humans more context and getting feedback for larger sections of behavior but also poses credit assignment problems. Finer-grained feedback is much easier to learn from and simplifies credit assignment, but is often impractical or tedious for humans to provide. It is also possible to strike a compromise, as demonstrated by Guan et al. (2021) who propose to use queries on step granularity, but batch queries within an episode to provide additional context and make it easier to give feedback. Note that we only classify a type of feedback as “episode” granularity if it requires entire episodes. If it is compatible with partial segments as well, we classify it as “segment” even if the discussed source paper uses entire episodes.
@@
This describes whether feedback is given directly on raw instances, e.g., behavior recordings (see granularity), or on abstract features of the instances, such as an autonomous car’s speed and the forces experienced by the passengers. While feature-level information can be easier to learn from, extracting useful features is challenging. In some contexts, it may also be harder for a human to make abstract judgments rather than more intuitive instance-level judgments. Note that this always refers to the level of abstraction that the user sees, which may differ from the features used as inputs for the reward model. Types of feedback that depend on active generation (see involvement), such as improvements, generally work on a raw instance level.
@@
Humans may communicate explicitly for the purposes of feedback, e.g., by responding to a pairwise comparison query, or implicitly as a side-effect of actions directed at other purposes such as interventions in the agent’s behavior. While explicit information is easiest to learn from, implicit information is often much more readily available and can possibly communicate more detailed preferences.
@@
The distinction between literal and pedagogical feedback was introduced by Milli & Dragan (2020). They argue that humans with pedagogical intent (i.e., evaluative, instructive, or descriptive) may act differently compared to humans with literal intent. Even though they find that assuming the (wrong) literal intent can still lead to better reward inference, it still indicates that it can be important to know this intent to choose the right human model (see Section˜5.1.5).
@@ 3.2 Common Classes
Table˜2 gives an overview of the classes and their attributes as described in Section˜3.1.
@@ 3.2.1 Primary Feedback Classes
Critique is arguably the most direct type of feedback. In this setting, the human expresses their preference by directly critiquing an instance of agent behavior, often in the form of binary feedback. Note that critique, as considered in this survey, is distinct from directly supplying a reward signal since it is given in an asynchronous and indirect manner (see Section˜1.3). In the critique setting, a labeler may for example observe recordings of agent behavior and give their approval or disapproval. The defining features of critique are that the human passively observes the behavior (involvement), gives feedback on a single instance (arity), and does so explicitly (explicitness) with an evaluative intent. The feedback may be given for any granularity and on any level of abstraction.
@@
Binary comparisons and rankings are among the most common types of feedback. The defining features of comparisons are that the human passively observes the behavior (involvement), gives relative feedback on multiple instances (arity), and does so explicitly (explicitness) with an evaluative intent. It is most commonly given on a segment (granularity), but other granularities such as individual actions are also possible (Cosner et al., 2022). Similarly, comparisons are commonly requested on an instance level (abstraction), but this is not a requirement.
@@
This basic setting has been extended and modified in various ways. To reduce noise in the labels, it is common to extend the binary choice by giving the labelers the option to avoid the hard choice and instead indicate incomparability, uncertainty, or perceived similarity (Holladay et al., 2016). This option is commonly interpreted as “equally preferable”, e.g., as if each trajectory had an equal probability of being preferred in the preference predictor. It is also common, however, to additionally provide an “incomparable” option which results in simply omitting the query from the data set (Christiano et al., 2017; Ibarz et al., 2018). In contrast to this, Verma et al. (2023) explicitly state that they do not allow for the equally preferred option, arguing that these cases are rare enough not to matter very much. Another line of research suggests that more precision in the expression of pairwise preferences, such as softening the hard binary choice to scalar feedback indicating the strength of a preference (Wilde et al., 2021; Touvron et al., 2023), can be beneficial for preference learning. Other extensions change the pairwise setting to choices among larger choice sets (Ziegler et al., 2020) or even full rankings (Myers et al., 2021; Ouyang et al., 2022). Jain et al. (2015) compare different kinds of re-ranking feedback (select first which is better than top, choose best from top-5, choose best from random set). Their study suggests that the first two are roughly equivalent, while the latter is much less informative. Askell et al. (2021) evaluate binary comparisons as well as ranking, but find that binary comparisons (even if extracted from a ranking) perform better. Ziegler et al. (2020) note that for language tasks, a larger choice set can amortize the time needed for a labeler to get acquainted with the context necessary to understand a query. Basu et al. (2019) propose to use hierarchical queries, i.e., a sequence of pairwise comparisons that build up on each other.
Pairwise comparisons are generally easier to supply than ratings, even when the feedback is given by a foundation model (Wang et al., 2024c). While absolute feedback is generally dependent on a policy serving as an implicit baseline (MacGlashan et al., 2017), the explicit baseline in pairwise comparisons can be more easily controlled. Therefore, comparisons provide many benefits over absolute ratings, such as reduced bias, inconsistencies, and subjectivity (Yannakakis & Martínez, 2015). On the flip-side, comparisons convey relatively little information per label. Tien et al. (2023) study the weaknesses of pairwise-comparison-based reward learning. Since the amount of information provided for each label is small, these models are prone to causal confusion, i.e., misattributing reward to noise and misidentifying the reward function.
@@
One limitation of trajectory comparisons is that they require a set of roughly comparable trajectories. In many real-world environments, starting conditions or even the agent’s current task may vary between episodes. In these cases, it is hard for human labelers to compare trajectories from these episodes, limiting the usefulness of comparison feedback. One way to remedy this limitation is to provide feedback within a single trajectory. Instead of comparing a set of instances with each other as in regular comparative feedback, inter-temporal feedback conveys relative judgments over different states in time within a single instance, e.g., by comparing a robot’s behavior at the start and the end of a trajectory. The defining features of inter-temporal feedback are that it is given explicitly (explicitness) on a segment (granularity) while passively observing (involvement) a single instance (arity) of the agent’s behavior with evaluative intent. It is most commonly given on raw instances, but any level of abstraction is possible in principle. There are two main ways to convey this feedback: Reward sketching and inter-temporal preferences.
Reward sketching, as introduced by Cabi et al. (2020), involves users sketching a visual representation of the reward function over time. This type of feedback, which can be given by sketching a graph with the mouse while watching a behavior recording, provides intuitive reward annotations for each time step. Rahtz et al. (2022) also adopted this approach, referring to it as “one of the highest-bandwidth feedback mechanisms currently available”.
@@
Proxy rewards are partial or inaccurate reward functions that convey information about the task the agent is supposed to complete but may not induce optimal behavior. An example of this is a reward function that only rewards the agent for reaching the goal state but does not penalize negative side-effects. This form of feedback does not generally refer to any particular behavior instance but instead gives global direction for the entire task. However, in line with our selection criteria (Section˜1.3), we only consider proxy reward feedback that is interactive and online within the context of one or multiple observations to fill holes in the initial description. The defining features of proxy reward feedback is that the labeler passively observes the agent’s behavior (involvement) and gives feedback explicitly (explicitness) on a feature-level (abstraction) with descriptive intent (intent). Proxy reward feedback may be given with respect to a single or multiple instances (arity), although it generally refers to multiple instances. It is most commonly given on an episode (granularity), but other granularities are possible in principle.
The work by He & Dragan (2021) exemplifies this form of feedback through an iterative process where designers provide proxy reward functions on training environments. The method maintains a belief distribution over reward functions consistent with behaviors that achieve high proxy reward on training environments, then actively proposes new environments that maximize information gain about the true reward function, helping designers refine their reward specifications. Alternatively, Mindermann et al. (2018) suggest querying about the reward function, extending the inverse reward design framework of Hadfield-Menell et al. (2017b), which treats proxy reward functions as observations to infer the designer’s true intent. The active approach by Mindermann et al. (2018) allows users to choose from a set of understandable, linear proxy rewards or to specify which features are more critical in the linear reward structure. In a related setting, Guan et al. (2023) lets the user specify changes to a current symbolic reward function in the form of changes to target attributes (e.g., walking speed).
@@
Interventions can be considered to be an instance of the shared autonomy setting since the agent and the user share autonomy to reach a common goal. There are two main ways to leverage interventions: One is to learn from the occurrence of an intervention itself that the prior behavior was suboptimal, as done, e.g., by Luo et al. (2024) by directly inferring negative rewards from interventions without learning a reward model, and the other is to learn from the corrected trajectory. The latter setting is studied by Abramson et al. (2022), who ask humans to intercede on agent failure. They use this to collect targeted demonstrations where the agent is the weakest and to identify challenging situations for their evaluations. The gathered data is then used for behavior cloning and reward model training. The fact that a correction occurred is not directly used as feedback. In addition to the above distinction, interventions can come in two main formats: Either by overtaking control from the agent to correct its behavior or by physically correcting the agent’s movements. Losey et al. (2022) proposes to learn a reward model from such physical corrections, a method that was later extended (Losey & O’Malley, 2018) to incorporate uncertainty for active learning and risk-sensitive deployment. Li et al. (2021) further extend this setting to learn from a sequence of correlated physical corrections without needing to wait until the trajectory is completed.
The correction case is closely related to the setting of coactive learning (Shivaswamy & Joachims, 2015), in which a learning system repeatedly proposes a solution which a user may correct to reach a common goal. Jain et al. (2015) treat corrections as a demonstration while Jeon et al. (2020) propose (but do not evaluate) an alternative interpretation of inferring implicit preferences from comparisons by assuming the corrected trajectory is preferred over the original one. Corrections are also commonly used in the LLM fine-tuning setting to incorporate AI feedback: By prompting a language model with a set of principles, it can critique and revise a generated response according to those principles. This can then be used to generate demonstrations (Bai et al., 2022b) or targeted pairwise preferences (Castricato et al., 2024) and to internalize these principles with further training on this data.
@@
While much work in language for RL focuses on the problem of learning a policy that follows natural-language instructions in a non-interactive setting (Hermann et al., 2017; Nair et al., 2021), we focus on works where language is used as feedback, not for task definition. Note, however, that RLHF can be a useful tool for learning a language-conditioned policy, regardless of whether or not the feedback itself is in the form of natural language. As an example of this, Abramson et al. (2022) use RLHF in an interactive agents setting, where one agent (responsible for solving problems) is controlled by the learned policy and another agent (responsible for setting tasks) is controlled by the human. RLHF for LLMs can also be viewed as language-conditioned RL, although Shen et al. (2024) find that reward models in this context often fail to distinguish nuances in task descriptions.
@@
Language feedback can also be interpreted by an LLM directly, without learning a reward model. This is demonstrated in the concurrent works by Ma et al. (2024b) and Xie et al. (2024) who propose to learn a symbolic reward model in the form of a language-model written piece of (Python) code. They use natural language feedback to improve the reward model based on observations of the agent’s behavior induced by the previous version of the reward model.
@@ 3.2.2 Supplementary Classes
This section introduces two feedback classes, e-stops and importance, that can be used in conjunction with a primary class to learn a reward model. These classes are not sufficient to learn a reward model on their own but can supplement a primary feedback type.
@@
Emergency stops (e-stops) (Ghosal et al., 2023) are an active type of feedback. In this type of feedback, the human may intervene with the agent’s behavior by stopping it, i.e., they may choose to stop the agent’s current trajectory at any point, for example to prevent the agent from causing harm. This is closely related to interventions but, in contrast to those, e-stops do not suggest an alternative action. The defining features of e-stops are that the human passively observes the agent’s behavior (involvement), gives absolute feedback on a single instance (arity) on the instance level (abstraction), and does so implicitly (explicitness) as a side-effect of regular interaction. The intent is literal due to the implicit nature. For the purposes of intervention, the human usually observes the full episode (granularity). Due to the small amount of infrequent information they provide, e-stops should only be considered as a supplementary feedback type.
E-stops are primarily intended to prevent bad behavior and only implicitly convey information about the correct behavior. This interaction and the arising incentives have been formalized in the form of the “off-switch game” by Hadfield-Menell et al. (2017a). Jeon et al. (2020) propose to interpret this as a form of reward-rational feedback, where the ‘off’ choice maps to the trajectory with the robot remaining still after the off switch has been triggered (see Section˜5.1.1). Kahn et al. (2021) demonstrate that a robot can learn to navigate using such feedback.
@@
Another form of supplementary feedback may come in the form of importance labels, communicating which parts of the observation are important for the objective, e.g., by marking salient regions in an image representing the agent’s final state or important words in the language model’s response. Its defining features are that the importance information itself does not contribute towards generating behavior samples (observed involvement), is of descriptive intent, and is given explicitly (explicitness). Granularity, arity, and abstraction may vary depending on the primary feedback type. Since importance feedback needs a base task with respect to which the importance is defined, it cannot be used on its own but is rather a supplementary type of feedback.